# Augmenting Math Word Problems via Iterative Question Composing

### 17 Jan 2024 v1

@@ Abstract
Despite recent progress in improving the mathematical reasoning ability of large language models(LLMs), solving competition-level math problems without the use of external tools remains challenging for open-source LLMs. In this work, we introduce the MMIQC dataset, a mixture of processed web data and synthetic question-response pairs, to equip base models with better mathematical reasoning skills. Mistral-7B-MMIQC, the model obtained by fine-tuning Mistral-7B(Jiang et al., 2023) on MMIQC, achieves 36.0% accuracy on MATH(Hendrycks et al., 2021b), 5.8% higher than the previous (model size $\sim$7B) SOTA. Our experiments also show that a large part of the improvement attributes to our novel augmentation method IQC(Iterative Question Composing), where we iteratively ask an LLM to compose new questions from the given seed problems and do rejection sampling from another LLM. We have released MMIQC on this HuggingFace hub page.
@@ 1 Introduction
We fine-tune some base models, resulting in models consistently achieve a large margin compared to their counterparts on MATH, as shown in Figure 1. Notably, Mistral-7B-MMIQC, the model obtained by fine-tuning Mistral-7B(Jiang et al., 2023) on MMIQC, achieves 36.0% accuracy on MATH, 5.8% higher than the previous (model size $\sim$7B) SOTA.
We emphasize our contributions as follows:
@@
  • We release MMIQC, a mixture of processed web data and synthetic question-response pairs. Notably, Mistral-7B-MMIQC, the model obtained by fine-tuning Mistral-7B(Jiang et al., 2023) on MMIQC, achieves 36.0% accuracy on MATH, 5.8% higher than the previous (model size smaller than 7B) SOTA.
@@ 4.1 Fine-tuning Setup
Our fine-tuning strategy mainly follows the practice of (Taori et al., 2023), except that we use a different prompt templater to transform the question-response pairs. For a sample from Mathematics Stack Exchange, the corresponding prompt fed into the model durning training is a simple concatenation of the question and response with two new-line symbols. For a sample from other subsets, we additionally add a prefix ‘Please solve the following problem and put your answer at the end with “The answer is: ”.’ to the question-response concatenation.
@@ 4.4 Contamination test
We check the $n$-gram matches for MMIQC to ensure that the improvement is not a result of direct memorization. We use the script provided by (Azerbayev et al., 2023) to check the $n$-gram matches between the synthetic part of MMIQC and MATH test set. It turns out that for a 30-gram match check, there are 44 hits of match between the ‘solution’ field of MATH test set and the ‘output’ field of MMIQC, far fewer than the 168 hits between that of MATH test set and MATH training set. Morever, we manually check these 44 hits and find that 43 among them are belong to the case where intermediate steps of the solutions to similar but different questions collide, with the only exception being the question ‘A regular polygon has interior angles of 144 degrees. How many sides does the polygon have?’. This almost rules out the possibility that fine-tuned models get memorization of solutions to the problems in the test set, indicating a very low risk of data contamination for MMIQC.

### 20 Jan 2024 v2

@@ Abstract
Despite recent progress in improving the mathematical reasoning ability of large language models(LLMs), solving competition-level math problems without the use of external tools remains challenging for open-source LLMs. In this work, we introduce the MMIQC dataset, a mixture of processed web data and synthetic question-response pairs, to equip base models with better mathematical reasoning skills. In different model sizes, the models fine-tuned on MMIQC consistently outperform their counterparts by a clear margin on MATH test set. Notably, DeepSeek-67B-MMIQC achieves a 41.0% accuracy, 4.2% higher than the previous open-source SOTA. Our experiments also show that a large part of the improvement can be attributed to our novel augmentation method IQC(Iterative Question Composing), where we iteratively ask an LLM to compose new questions from the given seed problems and do rejection sampling from another LLM. We have released MMIQC on this HuggingFace hub page.
@@ 1 Introduction
We fine-tune some base models on MMIQC, resulting in models consistently achieve a large margin compared to their counterparts when evaluated on MATH, as shown in Figure 1. Specifically, Mistral-7B-MMIQC obtained by fine-tuning Mistral-7B(Jiang et al., 2023) on MMIQC, Llemma-34B-MMIQC for Llemma-34B(Azerbayev et al., 2023) and DeepSeek-67B-MMIQC for (Bi et al., 2024) achieve 36.0%, 38.6% and 41.0% accuracy on MATH, 5.8%, 3.8% and 4.2% higher than the previous SOTAs in their model size levels, respectively.
We highlight our contributions as follows:
@@
  • We release MMIQC, a mixture of processed web data and synthetic question-response pairs. In different model sizes, the models fine-tuned on MMIQC consistently outperform their counterparts by a clear margin on MATH test set. Notably, DeepSeek-67B-MMIQC achieves a 41.0% accuracy, 4.2% higher than the previous open-source SOTA.
@@ 4 Experiments
DeepSeek-67B(Bi et al., 2024) N/A No Pass@1 18.7
Deepseek-67B-MetaMathQA MetaMathQA No pass@1 36.8
Deepseek-67B-MMIQC* MMIQC No pass@1 41.0
Deepseek-67B-MetaMathQA MetaMathQA No maj1@256 45.4
Deepseek-67B-MetaMathQA MetaMathQA Shepherd maj+verify1@256 48.1
@@ 4.1 Fine-tuning Setup
Our fine-tuning strategy mainly follows the practice of (Taori et al., 2023), except that we use a different prompt template to transform the question-response pairs. For a sample from Mathematics Stack Exchange, the corresponding prompt fed into the model during training is a simple concatenation of the question and response with two new-line symbols. For a sample from other subsets, we additionally add a prefix ‘Please solve the following problem and put your answer at the end with “The answer is: ”.’ to the question-response concatenation.
@@ 4.4 Contamination test
We check the $n$-gram matches for MMIQC to ensure that the improvement is not a result of direct memorization. We use the script provided by (Azerbayev et al., 2023) to check the $n$-gram matches between the synthetic part of the MMIQC and MATH test set. It turns out that for a 30-gram match check, there are 44 hits of match between the ‘solution’ field of MATH test set and the ‘output’ field of MMIQC, far fewer than the 168 hits between that of MATH test set and MATH training set. Moreover, we manually check these 44 hits and find that 43 among them belong to the case where intermediate steps of the solutions to similar but different questions collide, with the only exception being the question ‘A regular polygon has interior angles of 144 degrees. How many sides does the polygon have?’. This almost rules out the possibility that fine-tuned models get memorization of solutions to the problems in the test set, indicating a very low risk of data contamination for MMIQC.